OpenStudy (anonymous):

FAN AND MEDAL !! Let f(x) be even and g(x) be odd . Using the properties of even and odd functions , is f(g(x)) and g(f(x)) even, odd, or neither ?

OpenStudy (amistre64):

try an example ... f(x)=x^2 and g(x) = x

OpenStudy (kirbykirby):

An even function \(g\): means that \(g(−x)=g(x)\) An off function \(m\): means that \(m(−x)=−m(x)\) To test if the composition \(f(g(x))\) is even, you should check that \((f∘g)(−x)=(f∘g)(x)\) So: \((f∘g)(−x)=f(g(−x))=f(−g(x)) \)since \(g\) is odd. \(f(−g(x))=f(g(x))=(f∘g)(x) \) since\( f\) is even Can you try the same idea for the other composition?